A study is being done on the catfish population of a lake. An enviromental worker catches, tags, and frees 96 catfish. One month later, he catches and frees 116 catfish and 29 have tags. Which proportion should be used to estimate the total number of catfish, x, in the lake?
Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12569)

96/x=29/116
29x=96*116
29x=11,136
x=11,136/29
x=384 answer for the projected catfish population.

Answer by stanbon(48568)

A study is being done on the catfish population of a lake. An enviromental worker catches, tags, and frees 96 catfish. One month later, he catches and frees 116 catfish and 29 have tags. Which proportion should be used to estimate the total number of catfish, x, in the lake?
-----------------
(actual population)/(tagged and released) = (size of 2nd catch/tagged in 2nd catch)
----------------------
ap/96 = 116/29
ap = 96[116/29]
ap = 384
----------------
Cheers,
Stan H.